OpenStudy (anonymous):

lim x-0 (x+ sin4x)/tan3x+4x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x+\sin 4x }{ \tan 3 x+4x }\] \[=\lim_{x \rightarrow 0}\frac{ x+4x \frac{ \sin 4 x }{ 4 x } }{3 x \frac{ \tan 3 x }{ 3 x }+4 x }\] \[=\lim_{x \rightarrow 0}\frac{ x \left( 1+4\frac{ \sin 4 x }{ 4x } \right) }{ x \left( 3\frac{ \tan 3 x }{ 3x}+4 \right) }\] \[=\lim_{x \rightarrow 0}\frac{ 1+4 \frac{ \sin 4 x }{ 4 x } }{ 3 \frac{ \tan 3 x }{ 3x+4 } }\] \[=\frac{ 1+4*1 }{ 3*1+4 }=\frac{ 5 }{ 7 }\]